# LLM+KG’24: The International Workshop on Data Management Opportunities in Unifying Large Language Models + Knowledge Graphs, August 25, 2024, Guangzhou, China

### v1

@@ 2.5.2 Inconsistency Detection
This section delves into Research Question 3, examining the role of LLMs in identifying and fixing inconsistencies in KGs. KG consistency is about ensuring that the information in a KG is not self-contradictory[59]. KG consistency differs from KG accuracy in that it prioritizes the logical soundness of the KG rather than the factual correctness of individual triples. A KG might contain outdated yet logically coherent information, maintaining high consistency even with low accuracy. Conversely, high accuracy usually implies good consistency, as factual data tends to be logically harmonious[60]. Current rule-based techniques for KG completion and inconsistency detection primarily utilize structural information, overlooking the semantics of relations. Using LLMs can harness both the semantic and structural information from KGs to generate meaningful rules. Among recent works leveraging LLMs for rule-based methods, there is ChatRule [61], a framework that uses both the semantic and structural information of KGs to prompt LLMs to generate logical rules. This segment of the paper is dedicated to addressing Research Question 1, which investigates the effective use of LLMs in generating descriptive textual information for entities within KGs. KG-to-Text generation encompasses the procedure of transforming structured data held within a KG into human-readable, natural-language text [9]. Colas et al. combined a pre-trained LM with graph attention information by modifying the encoder layer. Initially, they convert the KG into a textual string by linearizing its structure. Then, the vector representation becomes contextualized through the incorporation of graph attention data. Several methods involve the fine-tuning of pre-trained language models specifically on KG-to-text datasets to address this particular task [10]. Chen et al. gathered a corpus named KGTEXT which involves matching linked sentences sourced from Wikipedia with knowledge subgraphs obtained from WikiData and then conducting pre-training and subsequently fine-tuning on KG-to-Text datasets. [12] developed a basic yet effective structure-aware semantic aggregation module that integrates into every Transformer layer within the encoder. This module’s purpose is to maintain the structural details present in the input graphs. Then followed by three pre-training tasks aimed at explicitly acquiring graph-text alignments. [13] investigated few-shot KG-to-text generation which generates the output with the help of around several hundred instances. Initially, their model aligns KG representations, encoded by graph neural networks, with PLM-based entity representations, aiming to close the semantic gap. It then proposes a specialized strategy, called relation-biased breadth-first search (RBFS), to arrange the KG into a well-structured entity sequence for integration into PLMs. Lastly, the model employs multi-task learning, simultaneously training the main text generation and an auxiliary KG reconstruction task. This approach significantly improves the semantic alignment between the input KG and the resulting text, allowing the model to generate accurate descriptions of the KG.
Ontology Creation: Concept Extraction to Validation: This segment of the paper specifically addresses Research Question 2, delving into the application of LLMs in ontology generation for KGs. LLMs exhibit significant promise in the construction of ontologies [14], where the organized representation of knowledge is essential for facilitating intelligent information retrieval and reasoning. LLMs contribute to creating and enriching comprehensive ontologies, with domain-specific fine-tuning enhancing their integration into specialized areas. LLMs are promising for ontology generation but face computational intensity and bias challenges, requiring broader evaluation, yet remaining valuable in ontology creation and intelligent systems.
Concept and Relation Extraction: LLMs like GPT-3, with their natural language understanding abilities, are suitable for extracting ontology concepts and relations. Concept extraction is the process of identifying and extracting concepts from text. The process involves extracting linguistic realizations of domain-specific concepts and accumulating semantic term variations with sense disambiguation and domain-specific synonym identification [16]. Relation extraction aims to identify non-taxonomic relations between concepts [17]. It aims to identify relations between text entities and can be enhanced by using techniques like Named Entity Recognition (NER) and additional context provided by LLMs [18, 19].
Property Identification: Research investigates using fine-tuned LLMs for pre-annotating data in ontology property identification, helping annotators efficiently discern relevant properties and their correlation with automated scores, thereby reducing annotation times.[20].
Ontology Enrichment: LLMs can help to improve current ontologies by adding new information or improving concept definitions. This is especially beneficial when working with dynamic domains in which the ontology must adapt to new facts. Advanced machine learning methods, particularly fine-tuned LLMs, have been researched for data pre-annotation and ontology enrichment [20].
Ontology Alignment: LLMs can help to align ontologies by understanding and comparing the semantics of ideas and relationships among ontologies. Ontology alignment with LLMs has been studied in the literature. One method suggests a neurosymbolic design that combines the flexibility of LLMs and the domain orientation of Enterprise KGs (EKGs) [21]. Another work investigates the use of fine-tuned LLMs to pre-annotate data in a lexical extension job, specifically adding descriptive words to an existing ontology [20].
Text-to-Ontology Mapping: Train LLMs to map natural language text to specific ontology concepts and relationships. This can be accomplished using supervised learning, which involves training the model using annotated instances of text-to-ontology mappings. Text-to-ontology mapping using LLMs has been explored in multiple papers. One approach involves using artificial neural networks and classifiers to match texts to relevant ontologies [22].
Ontology Learning automates the construction of ontologies by identifying concepts from text, aiming for an effective, scalable knowledge representation across disciplines. Recent advances in LLMs like BERT offer a promising alternative to traditional, time-consuming ontology creation, outperforming classic word embedding models in various nlp tasks and providing a more efficient method for knowledge acquisition [16, 23, 17, 24].
2.5.3 Entity Extraction and Alignment:
Named Entity Recognition (NER) task involves being provided with a sentence, and the objective is to predict the set of entities present within that sentence[25]. Some papers concentrated on prompting to extract entities. [25] approach named PromptNER, comprises four essential components: a backbone LLM, a prompt outlining the set of entity types, a small set of examples from the target domain, and a specific format for outputting the extracted entities. Some approaches employ instruction-tuning. It is a technique in which pre-trained autoregressive language models undergo finetuning to adhere to natural language instructions and produce responses [26]. Zhou et al. proposed a focused distillation method with instruction tuning tailored to a specific mission.
2.5.4 Relation Extraction
LLMs have notably advanced the complex task of relation extraction between text entities, revolutionizing the identification and interpretation of semantic connections in textual data. In particular, GPT-RE [28] introduced a novel approach by infusing task-specific knowledge and logic-enriched demonstrations. This approach effectively reduces the low relevance issues between entity-relation pairings and overcomes the limitations in explaining input-label correlations. Furthermore, considering the vast array of predefined relation types and the challenges posed by uncontrolled LLMs, the study [29] suggests incorporating a natural language inference module within LLMs. This integration focuses on generating relation triples to greatly improve the usefulness and precision of document-level relation datasets. These advancements in relation extraction demonstrate LLMs’ capability to improve semantic analysis in text, and we have organized these methods by their distinct learning approaches. This classification comprises:
Supervised Fine-tuning known for being prevalent and promising, involves extensively inputting training data to fine-tune LLMs. It allows LLMs to learn and comprehend the structural patterns in data, enhancing their ability to generalize. The Deepstruct model, referenced by [30], uses structure pre-training on diverse corpora to improve language models’ understanding of complex language structures, leading to a more nuanced interpretation of linguistic patterns. [31] enhanced relation extraction by training a specialized BART-style model [32] with a novel triplet linearization technique. This development demonstrates the importance of supervised fine-tuning in advancing LLMs for complex tasks like relation extraction.
Few-shot learning, which relies on a limited number of labeled examples, faces significant challenges in machine learning. The lack of data can cause problems like overfitting and difficulty in understanding complex data relationships, as pointed out by [33]. The advent of LLMs has been a significant breakthrough in overcoming few-shot learning challenges. [34] introduced two strategies for using LLMs in relation extraction: 1) in-context learning (ICL) and 2) data generation. ICL involves using detailed prompts with task definitions and labels to help LLMs understand relation extraction nuances, particularly useful in few-shot settings with limited data. To further address data scarcity, they used LLMs to generate additional training material by creating prompts with instance descriptions and examples. Complementarily, [35] examined the effectiveness of LLMs in few-shot relation extraction through in-context learning. They discovered that LLMs could equal state-of-the-art techniques using just a few examples. Their method included enhancing relation extraction labels with Chain of Thought (CoT) explanations from GPT-3 and fine-tuning the Flan-T5 model with this enriched data.
Zero-shot learning aims to empower models to generalize to new tasks and domains beyond their training, by adapting the pre-training of LLMs to unfamiliar situations. The extensive knowledge embedded in LLMs demonstrates their remarkable capability for zero-shot tasks in various fields. Key studies such as those by [36, 37], and recent advancements [38], exemplify the potential of LLMs in adapting to tasks they were not explicitly trained for. Adding to this discourse, [39] critically assessed ChatGPT’s zero-shot learning capability in temporal relation extraction. The study emphasized ChatGPT’s skill in grasping complex temporal relations but also noted its limitations in consistency and handling long-dependency relations.
2.6 KG Reasoning
@@
Recently, there has been a growing emphasis on the exploration of KG reasoning by the potential of this approach to draw new knowledge and conclusions from the wealth of existing data [40]. It has applications for KG completion and KG question answering[40]. Luo et al. introduced Reasoning on Graphs(RoG), which combines LLMs and KGs to perform reliable and easily understandable reasoning. To tackle challenges such as hallucinations and knowledge deficiencies, they proposed a framework for planning-retrieval-reasoning. Initially, the planning module produces relation paths in the form of faithful plans using KGs, forming the basis for the RoG approach. Subsequently, these devised plans are employed to extract legitimate paths of reasoning from KGs to carry out reliable reasoning through the retrieval-reasoning module.
2.7 KG Completion
@@
In KG completion, the main tasks are: (i) triple classification, validating the accuracy of a given triple, (ii) link prediction for predicting missing components like the head, tail, or relation in a triple, and (iii) entity classification for categorizing entities, also known as determining the entity’s type[42]. The triple-based methods relying on structural information, are widely used for KG completion, mostly the embedding methods [43, 44, 45]. Several works have attempted to prove more effectiveness than triple-based completion by relying on the extraction of factual knowledge from LMs embedding spaces, instead of simple structures. Examples of these text-based methods, KG-BERT [46], and the fine-tuning of GPT-2 [47]. KG-BERT[46] is a leading example of applying LLMs to KG completion. It fine-tunes for this task by treating triples as textual sequences. The model uses KG-BERT to score sequences of entity-relation-entity, representing these elements through their names or descriptions. These sequences of words are then used as input for fine-tuning the BERT model. KG-BERT was identified with two shortcomings: 1) the model’s struggle to learn relational data in KGs, and 2) difficulty in identifying the correct answers among similar candidates. To improve this, multi-task learning methods combining relation prediction and relevance ranking with link prediction were proposed, as in [48]. This enhanced model better grasps KG relationships and effectively handles lexical similarities. In another study, [49]introduced the Structure-augmented Text Representation (StAR) model, aimed at improving link prediction in KG completion. The model begins with a Siamese-style textual encoder applied to triples, generating two contextualized representations.A distinctive scoring module using dual strategies captures both contextualized and structured knowledge, enhanced by a self-adaptive ensemble scheme with graph embedding techniques to boost performance. [50] presented SimKGC, a straightforward approach to enhance text-based KG completion. It recognizes the crucial aspect of conducting efficient contrastive learning. Building on recent advancements in contrastive learning. SimKGC utilizes a bi-encoder architecture and incorporates three types of negative examples to achieve its objectives. Sequence-to-sequence PLM methods have also shown state-of-the-art performance by redefining the KG completion challenge as a sequence-to-sequence generation task as demonstrated in KG-S2S [KG_S2S] and GenKGC [52]. GenKGC incorporates innovative techniques such as relation-guided demonstration and entity-aware hierarchical decoding, aimed at improving the quality of representation learning and accelerating the inference process. Most of these recent text-based methods require specific fine-tuning for KGs, which limits their efficiency. To address this limitation, new training free methods with no fine-tuning were recently proposed, as in KICGPT [53].
2.8 KG Embedding
@@
KG Embedding involves mapping elements of a KG into vectors in a continuous vector space. Unlike GNNs, which learn KG embeddings from data explicitly represented by nodes and edges, LM learns KG embeddings from natural language. Given a piece of text, LMs learn representations of entities, which capture context information containing relations of entities. The entity relations are found by an attention mechanism, which can create an association between two entities independent of their location. The more text an LM sees the more entities and their relationships the LM can learn; the larger an LM the more contextual information the representation contains. LLMs contain billions of parameters and are trained over a huge amount of textual data (e.g. 45 TB for GPT-3). LLMs perfectly meet these requirements and so are a powerful tool for learning KG embeddings. LLMs’ ability to KG embedding attracts researchers to use them directly to conduct KG tasks, like entity alignment [54]. We can also use the representation of entities learned by LLMs in the small-sized models. This should significantly reduce the amount of training data needed and the time of training and create well-generalized models. An extensive experiment is needed to investigate the efficiency of applying embeddings of LLMs into small-sized models for KG analysis tasks.
2.9 KG Validation
@@
Fact Checking: In this section, we address Research Question 4, which is centered on how LLMs can enhance the accuracy, consistency, and completeness of KGs through the process of fact-checking. The advent of LLMs has the potential to significantly impact the fight against misinformation, presenting both opportunities and challenges. LLMs, with their extensive world knowledge and strong reasoning abilities, offer promise in combating misinformation. However, the flip side is that these models can also be easily manipulated to generate deceptive misinformation at scale [55]. An approach to address misinformation in a KG involves verbalizing each triple within the KG and prompting LLMs to identify instances of misinformation [56, 1]. To address the limitations of potentially outdated or insufficient knowledge in LLMs for detecting factual errors, some research has explored augmenting LLMs with external knowledge[57] or utilizing tools [58] for misinformation detection. The methodologies developed for factual checking in LLMs can be applied to KG fact-checking by converting KG triples into textual representations.
@@
Inconsistency Detection: This section delves into Research Question 3, examining the role of LLMs in identifying and fixing inconsistencies in KGs. KG consistency is about ensuring that the information in a KG is not self-contradictory[59]. KG consistency differs from KG accuracy in that it prioritizes the logical soundness of the KG rather than the factual correctness of individual triples. A KG might contain outdated yet logically coherent information, maintaining high consistency even with low accuracy. Conversely, high accuracy usually implies good consistency, as factual data tends to be logically harmonious[60]. Current rule-based techniques for KG completion and inconsistency detection primarily utilize structural information, overlooking the semantics of relations. Using LLMs can harness both the semantic and structural information from KGs to generate meaningful rules. Among recent works leveraging LLMs for rule-based methods, there is ChatRule [61], a framework that uses both the semantic and structural information of KGs to prompt LLMs to generate logical rules.
3 KG-enhanced LLM
K-BERT [62] proposed to inject relevant domain knowledge from a KG into a sentence using a knowledge layer before the embeddings are calculated. The authors show that it improves performance in many NLP tasks. Sem-K-BERT [3] further extended this by integrating semantic role labeling (SRL) and by introducing semantic correlation calculation to reduce the noise. KCF-NET [63] demonstrated that incorporating KG representations with context helped to improve the LM performance in machine reading comprehension. Ji et al. proposes a concept-enhanced pertaining approach by adding concept semantics to setentences from external sources such as KGs. Dict-BERT [65] leverages definitions of the rare words from dictionaries to improve performance of various NLP tasks.
4 LLM-KG Cooperation
4.1 KG Question Answering

### v2

@@ 2.2. KG-to-Text Generation:
This segment of the paper is dedicated to addressing Research Question 1, which investigates the effective use of LLMs in generating descriptive textual information for entities within KGs. KG-to-Text generation encompasses the procedure of transforming structured data held within a KG into human-readable, natural-language text (Colas et al., 2022). Colas et al. combined a pre-trained LM with graph attention information by modifying the encoder layer. Initially, they convert the KG into a textual string by linearizing its structure. Then, the vector representation becomes contextualized through the incorporation of graph attention data. Several methods involve the fine-tuning of pre-trained language models(PLM), specifically on KG-to-text datasets, to address this particular task (Ribeiro et al., 2020). Chen et al. gathered a corpus named KGTEXT, which involves matching linked sentences sourced from Wikipedia with knowledge subgraphs obtained from WikiData and then conducting pre-training and subsequently fine-tuning on KG-to-Text datasets. Ke et al. developed a basic yet effective structure-aware semantic aggregation module that integrates into every Transformer layer within the encoder. The purpose of this module is to maintain the structural details present in the input graphs. Then followed by three pre-training tasks aimed at explicitly acquiring graph-text alignments. Li et al. investigated few-shot KG-to-text generation, which generates the output with the help of around several hundred instances. Initially, their model aligns KG representations, encoded by graph neural networks, PLM-based entity representations, aiming to close the semantic gap. It then proposes a specialized strategy, called relation-biased breadth-first search (RBFS), to arrange the KG into a well-structured entity sequence for integration into PLMs. Lastly, the model employs multi-task learning, simultaneously training the main text generation and an auxiliary KG reconstruction task. This approach significantly improves the semantic alignment between the input KG and the resulting text, allowing the model to generate accurate descriptions of the KG.
2.3. KG Reasoning
@@
KG Reasoning involves deriving new facts from existing ones using logical inference. Recently, there has been a growing emphasis on exploring KG reasoning due to its potential to derive new knowledge and conclusions from existing data(Chen et al., 2020a). It has applications for KG completion and KG question answering (Chen et al., 2020a). Choudhary and Reddy developed Language-guided Abstract Reasoning over Knowledge graphs (LARK) model that is designed to harness the reasoning power of LLMs for efficiently answering first-order logic (FOL) queries on knowledge graphs. LARK identifies relevant subgraph contexts within abstract KGs by utilizing the entities and relationships in the queries. It then conducts chain reasoning over these contexts using prompts from LLMs to break down and process logical queries. Luo et al. introduced Reasoning on Graphs (RoG), which combines LLMs and KGs to perform reliable and easily understandable reasoning. They proposed a framework for planning-retrieval-reasoning to tackle challenges such as hallucinations and knowledge deficiencies. Initially, the planning module produces relation paths in the form of faithful plans using KGs, forming the basis for the RoG approach. Subsequently, these devised plans are employed to extract legitimate paths of reasoning from KGs to carry out reliable reasoning through the retrieval-reasoning module. Kim et al. introduced KG-GPT, a versatile framework for reasoning on KG that operates in three stages: Sentence Segmentation, Graph Retrieval, and Inference. These stages are designed to break down sentences, extract relevant components from the graph, and draw logical conclusions.
2.4. KG Completion
@@
In KG completion, the main tasks are (i) triple classification, validating the accuracy of a given triple; (ii) link prediction for predicting missing components like the head, tail, or relation in a triple. And (iii) entity classification for categorizing entities, also known as determining the entity’s type (Alam et al., 2023). The triple-based methods relying on structural information are widely used for KG completion, mostly the embedding methods (Bordes et al., 2013; Lin et al., 2015; Trouillon et al., 2016). Several works have attempted to prove more effectiveness than triple-based completion by extracting factual knowledge from LMs embedding spaces instead of simple structures. Examples of these text-based methods are KG-BERT (Yao et al., 2019) and the fine-tuning of GPT-2 (Biswas et al., 2021). KG-BERT (Yao et al., 2019) is a leading example of applying LLMs to KG completion. It fine-tunes for this task by treating triples as textual sequences. The model uses KG-BERT to score sequences of entity-relation-entity, representing these elements through their names or descriptions. These sequences of words are then used as input for fine-tuning the BERT model. KG-BERT was identified with two shortcomings: 1) the model’s struggle to learn relational data in KGs, and 2) difficulty in identifying the correct answers among similar candidates. To improve this, multi-task learning methods combining relation prediction and relevance ranking with link prediction were proposed, as in (Kim et al., 2020). This enhanced model better grasps KG relationships and effectively handles lexical similarities. In another study, (Wang et al., 2021b) introduced the Structure-augmented Text Representation (StAR) model aimed at improving link prediction in KG completion. The model begins with a Siamese-style textual encoder applied to triples, generating two contextualized representations. A distinctive scoring module using dual strategies captures both contextualized and structured knowledge, enhanced by a self-adaptive ensemble scheme with graph embedding techniques to boost performance. Wang et al. presented SimKGC, a straightforward approach to enhance text-based KG completion. It recognizes the crucial aspect of conducting efficient contrastive learning. Building on recent advancements in contrastive learning. SimKGC utilizes a bi-encoder architecture and incorporates three types of negative examples to achieve its objectives. Sequence-to-sequence PLM methods have also shown state-of-the-art performance by redefining the KG completion challenge as a sequence-to-sequence generation task as demonstrated in KG-S2S (Chen et al., 2022) and GenKGC (Xie et al., 2022). GenKGC incorporates innovative techniques, such as relation-guided demonstration and entity-aware hierarchical decoding, aimed at improving the quality of representation learning and accelerating the inference process. Most of these recent text-based methods require specific fine-tuning for KGs, which limits their efficiency. New training-free methods with no fine-tuning were recently proposed to address this limitation, as in KICGPT (Wei et al., 2023b).
2.5. KG Embedding
@@
KG Embedding involves mapping elements of a KG into vectors in a continuous vector space. Unlike Graph Neural Networks, which learn KG embeddings from data explicitly represented by nodes and edges, LM learns KG embeddings from natural language. Given a piece of text, LMs learn representations of entities, which capture context information containing relations of entities. The entity relations are found by an attention mechanism, which can create an association between two entities independent of their location. The more text an LM sees the more entities and their relationships the LM can learn; the larger an LM, the more contextual information the representation contains. LLMs contain billions of parameters and are trained over a huge amount of textual data (e.g. 45 TB for GPT-3). LLMs perfectly meet these requirements, and so are a powerful tool for learning KG embeddings. LLMs’ ability to KG embedding attracts researchers to use them directly to conduct KG tasks, like entity alignment (Lippolis et al., 2023). We can also use the representation of entities learned by LLMs in the small-sized models, and this should significantly reduce the amount of training data needed and the time of training and, meanwhile, create well-generalized models. An extensive experiment is needed to investigate the efficiency of applying embeddings of LLMs into small-sized models for KG analysis tasks.
2.6. KG Validation
2.6.1. Fact Checking
@@
In this section, we address Research Question 4, which is centered on how LLMs can enhance the accuracy, consistency, and completeness of KGs through the process of fact-checking. The advent of LLMs has the potential to significantly impact the fight against misinformation, presenting both opportunities and challenges. LLMs, with their extensive world knowledge and strong reasoning abilities, offer promise in combating misinformation. However, the flip side is that these models can also be easily manipulated to generate deceptive misinformation at scale (Chen and Shu, 2023b). An approach to address misinformation in a KG involves verbalizing each triple within the KG and prompting LLMs to identify instances of misinformation (Chen and Shu, 2023a; Bang et al., 2023). To address the limitations of potentially outdated or insufficient knowledge in LLMs for detecting factual errors, some research has explored augmenting LLMs with external knowledge (Cheung and Lam, 2023) or utilizing tools (Chern et al., 2023) for misinformation detection. The methodologies developed for factual checking in LLMs can be applied to KG fact-checking by converting KG triples into textual representations.
2.6.2. Inconsistency Detection
@@
This section delves into Research Question 3, examining the role of LLMs in identifying and fixing inconsistencies in KGs. KG consistency is about ensuring that the information in a KG is not self-contradictory (Zaveri et al., 2016). KG consistency differs from KG accuracy in that it prioritizes the logical soundness of the KG rather than the factual correctness of individual triples. A KG might contain outdated yet logically coherent information, maintaining high consistency even with low accuracy. Conversely, high accuracy usually implies good consistency, as factual data tends to be logically harmonious (Wang et al., 2021a). Current rule-based techniques for KG completion and inconsistency detection primarily utilize structural information, overlooking the semantics of relations. Using LLMs can harness both the semantic and structural information from KGs to generate meaningful rules. Among recent works leveraging LLMs for rule-based methods, there is ChatRule (Luo et al., 2023a), a framework that uses both the semantic and structural information of KGs to prompt LLMs to generate logical rules.
3. KG-enhanced LLM
K-BERT (Liu et al., 2020) proposed to inject relevant domain knowledge from a KG into a sentence using a knowledge layer before the embeddings are calculated. The authors show that it improves performance in many NLP tasks. Sem-K-BERT (Xu et al., 2021) further extended this by integrating semantic role labeling (SRL) and by introducing semantic correlation calculation to reduce the noise. KCF-NET (Gong et al., 2020) demonstrated that incorporating KG representations with context helped to improve the LM performance in machine reading comprehension. Ji et al. proposes a concept-enhanced pertaining approach by adding concept semantics to sentences from external sources such as KGs. Dict-BERT (Yu et al., 2022) leverages definitions of rare words from dictionaries to improve the performance of various NLP tasks. Retrieval Augmented Generation (RAG) was developed to address issues such as hallucination, insufficient domain-specific knowledge, and outdated information by incorporating relevant external knowledge into the LLM prompt (Gao et al., 2023). RAG has three steps: Retrieval, Generation, and Augmentation. Over time, RAG has developed into three distinct models: Naive RAG, Advanced RAG, and Modular RAG. Naive RAG operates through a three-step process: indexing, retrieval, and generation. During the indexing step, text data is divided into small segments, each of which is encoded into vector form using embedding models. In the retrieval phase, the user’s query is similarly converted into a vector. The system then computes the similarity between the query vector and the vectors of the text segments, selecting the top k segments as the context for the response. Finally, in the generation step, both the query and the chosen contexts are input into an LLM as a prompt to generate a response. The Modular RAG has the ability to retrieve pertinent information from knowledge graphs(Gao et al., 2023). (Wang et al., 2023b) introduced KnowledgeGPT, a system that enhances LLMs with data from knowledge bases. To respond to a user query, their method generates a search code and executes it to retrieve information from the knowledge base. The output is then given to the LLM to respond to the user query. While RAG is effective in responding to queries directly related to specific excerpts from the source data where the retrieval is explicit, it may struggle with more general queries. For example, it might not perform well when asked, ”What are the main points of the dataset?” To overcome this limitation, Graph RAG was introduced (Edge et al., 2024). This approach generates a summary over a knowledge graph that has been constructed by the LLM from the source data.
Figure 2. Statistics of the usage of LLMs and KGs in cited papers per category
4. LLM-KG Cooperation
4.1. KG Question Answering
This section covers research questions 5 and 6.
4.1.1. Multi-Hop Question Generation: